OpenStudy (like1234):

What is the mean of this set of numbers? 7, –9, 6, –10, –13, –5 A. 4 B. –4 C. 7.7 D. –24

OpenStudy (askmelots):

hello im here to help :)

OpenStudy (like1234):

k

OpenStudy (like1234):

i think teh answer is B

OpenStudy (anonymous):

like1234 i think you need to sudtract those numberes

OpenStudy (like1234):

k

OpenStudy (askmelots):

The mean is the average of the numbers. It is easy to calculate: add up all the numbers, then divide by how many numbers there are. In other words it is the sum divided by the count. now I can help! XD

OpenStudy (like1234):

ok thanks thats what i did i just wanted someone 2 check my answer

OpenStudy (askmelots):

7 + -9 + 6 + -10 + -13 + -5=-24 -24/6=-4

OpenStudy (askmelots):

I got 6 from how many numbers there was

OpenStudy (askmelots):

your answer is -4

OpenStudy (anonymous):

the answer is c '

OpenStudy (askmelots):

Its B. -4

OpenStudy (anonymous):

ok b

OpenStudy (askmelots):

lol

OpenStudy (like1234):

thx guys!